OpenStudy (anonymous):

The mechanics at Lincoln Automotive are reboring a 6-in deep cylinder to fit a new piston. The machine they are using increases the cylinders radius one-thousandth of an inch every 4min. How rapidly is the cylinder volume increasing when the bore (diameter) is 3.800in

OpenStudy (anonymous):

okay, you can write it down like this in every 4 mins ---> radius increase = 1000in in every x mins? ----> radius increase R= D/2 =1.504 after finding the answer in mins, convert to seconds :) give it a try now :)

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

wait, I misread the question

OpenStudy (anonymous):

they want how much is the volume increasing right?

OpenStudy (anonymous):

Its okay :)

OpenStudy (anonymous):

at what rate is the volume increasing, supposed to use derivitives (this is calculus)

OpenStudy (anonymous):

then find the derivative of Volumeof Cylinder

OpenStudy (anonymous):

you have r' = 1000

OpenStudy (anonymous):

all you have to do is find h from the original equation which is Vol = 1/3 pi r^2 h using r = 1.504

OpenStudy (anonymous):

\[Vol = 1/3 \pi r^2h , Vol' = 2/3 \pi rr'h'\] I guess

OpenStudy (anonymous):

you have r' = 1000 r= 1.504 you'll have to find h'

OpenStudy (anonymous):

am I making any sense lol?

OpenStudy (anonymous):

Not really...

OpenStudy (anonymous):

How did you get r' ?

OpenStudy (anonymous):

jmcwhs2012, if you allow, may I say a few words

OpenStudy (anonymous):

it's given, it says that the radius is increasing 1000 in every 4 mins so r' = 1000in/min

OpenStudy (anonymous):

iam please go ahead :)

OpenStudy (anonymous):

sure:)

OpenStudy (anonymous):

"one-thousandth of an inch"

OpenStudy (anonymous):

Okay, as far as I understand, the rate at which the radius is increasing is equal to the radius bored in a second

OpenStudy (anonymous):

you have the original radius, you can computer r' I guess it's 1000(1.504)

OpenStudy (anonymous):

So we can write dr/dt= (2.54/100000)/(60*4)

OpenStudy (anonymous):

compute*

OpenStudy (anonymous):

Now all we need to find is dv/dt

OpenStudy (anonymous):

I dont understand where you got dr/dt?

OpenStudy (anonymous):

So we can use the result (dr/dt)x(dv/dr)

OpenStudy (anonymous):

(dr/dt)x(dv/dr)=dv/dt

OpenStudy (anonymous):

@jmcwhs2012 the thing is coming like this 2.54 cm is equal to one inch,

OpenStudy (anonymous):

Why do we have to use cm?

OpenStudy (anonymous):

I later convert it into m

OpenStudy (anonymous):

You see for that I divide it by 100000

OpenStudy (anonymous):

Ok let me explain the whole thing

OpenStudy (anonymous):

The machine is working at the rate of 1/1000 inch per 60*4 secs

OpenStudy (anonymous):

is that clear?

OpenStudy (anonymous):

wait @iam , he wants to find V'

OpenStudy (anonymous):

Thats what I am doing, if you all allow me

OpenStudy (anonymous):

lol, sorry, proceed :)

OpenStudy (anonymous):

The machine is working at the rate of 1/1000 inch per 60*4 secs, is that thing clear?

OpenStudy (anonymous):

yes, got you.

OpenStudy (anonymous):

Which means 2.54/100000 m per 240 secs

OpenStudy (anonymous):

So now we can find dr/dt from there

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

But we need to find dv/dt

OpenStudy (anonymous):

so, let's find it

OpenStudy (anonymous):

So we can use the formula dv/dt=(dr/dt)x(dv/dr)

OpenStudy (anonymous):

Is this step clear

OpenStudy (anonymous):

honestly, no, why are you multiplying them?

OpenStudy (anonymous):

wait! chain rule?

OpenStudy (anonymous):

Write that thing on a piece of paper, and try to cancel the commong terms, you will get why I am doing that

OpenStudy (anonymous):

Yes chain rule

OpenStudy (anonymous):

lol, got you! okay then what?

OpenStudy (anonymous):

So we have already found dr/dt which is 2.54/100000/240

OpenStudy (anonymous):

Now we need to find dv/dr

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

For that we will have to differenciate 6 x pi x r^2 with respect to r

OpenStudy (anonymous):

why 6 pi r^2?

OpenStudy (anonymous):

Oops I am sorry, it would be 6*2.54*pi*r^2

OpenStudy (anonymous):

6 is the length of the cylinder given on the problem

OpenStudy (anonymous):

I don't get where you got 6?

OpenStudy (anonymous):

oh, the height

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So now just we will have to multiply both the things

OpenStudy (anonymous):

but isn't the vol = 1/3 pi r^2 h?

OpenStudy (anonymous):

Thats for a cone

OpenStudy (anonymous):

which will be 2 pi r^2?

OpenStudy (anonymous):

lol...sorry ^^"

OpenStudy (anonymous):

So we just need to multiply both the things

OpenStudy (anonymous):

then differentiate with respect to r :)

OpenStudy (anonymous):

But unfortunately the person who gave this problem is not much interested

OpenStudy (anonymous):

No no no need to differenciate again

OpenStudy (anonymous):

lol, he'll come back later :)

OpenStudy (anonymous):

Its over

OpenStudy (anonymous):

so V = 15.34 pi r^2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what is r? 1.504?

OpenStudy (anonymous):

I didn't get you

OpenStudy (anonymous):

r is a variable

OpenStudy (anonymous):

wouldn't you derive to V' then substitute r = 1.504 and r' in the equation and then u are done?

OpenStudy (anonymous):

No not 1.504, we will have to substitute 1.900

OpenStudy (anonymous):

then by then you have found V' which is the volume rate

OpenStudy (anonymous):

I didn't get you

OpenStudy (anonymous):

alright, 1.900 and r' the one you have found which is dr/dt, then you're done

OpenStudy (anonymous):

right? :)

OpenStudy (anonymous):

Actually I am not getting you at all

OpenStudy (anonymous):

{(2.54/100000)/240}*2.54*6*pi*2*1.900=dv/dt Thats it and thats all

OpenStudy (anonymous):

so the given is: r = 1.900 r' = dr/dt = 2.54 so V' = 15.34(2) pi rr' = 30.68 pi (1.900)(2.54) = 465.15 m / secs So the volume is increasing = 465.15m/sec right?

OpenStudy (anonymous):

simply ^_^

OpenStudy (anonymous):

dr/dt = 2.54 this is not right

OpenStudy (anonymous):

so then it's 2.54 /1000/240?

OpenStudy (anonymous):

= 1.05 x10^-5

OpenStudy (anonymous):

dr/dt ={(2.54/100000)/240} dv/dr= 2.54*6*pi*2*1.900

OpenStudy (anonymous):

so: V' = dv/dr = 2.54*12 pi *1.900 = 181.94 m/secs